OpenStudy (princess123):

A rectangular park is three blocks longer than it is wide. The area of the park is 40 square blocks. If w represents the width, write an equation in terms of w for the area of the park and use it to find the length and width of the park. length= ? blocks width=? blocks

OpenStudy (anonymous):

\[L=w+3\] \[LW=40\] \[(W+3)W=40\] \[W^2+3 W=40\] \[W^2+3W-40=0\] \[(W-5)(W+8)=40\] \[W=5\]or \[W=-8\] so W=5

OpenStudy (princess123):

and sorry what is the length is the width is 5?

OpenStudy (gw2011):

W = Width L = Length = W+3 A = (L)(W) 40 = (W+3)(W) 40 = W^2+3W W^2+3W-40 = 0 (W+8)(W-5) = 0 W = -8 W = 5 Width must equal 5 and not -8 since a rectangle can't have a negative measurement L = W+3 = 5+3 = 8 Therefore, W = Width = 5 and L = Length = 8